OpenStudy (anonymous):

find all values of x for which the tangent to y=(x+2)^2 passes through the origin

OpenStudy (dumbcow):

The slope at any point is dy/dx dy/dx = 2x + 4 Equate this slope with the slope between the origin (0,0) and any given point (x,y) slope = (y2-y1)/(x2-x1) = y-0/x-0 = (x+2)^2/x --> (x+2)^2/x = 2x+4 (x+2)^2 = x(2x+4) x^2 +4x +4 = 2x^2 +4x x^2 = 4 x = +-2 At